3. [6 marks] Solve the IVP: cos x + y sin x = sinº x, y(0) = 2.
Answer #1

sol Given differential equations is cosx. dys & y. sinxe- Sinek) ; y(o)=2 Rewrite it as: of ty taux = sunt = d+yo tause = sinSo sol is given by y(x). IF = /QW. D. F du + C = y(x). Secx= /sinsritann. Secxdx +C = Staunda + c = f(secx-1) dx + c = taux-